OpenStudy (anonymous):

3/5 of the voters promised to vote for X and the rest promised to vote for Y. Of these, on the last day, 15% of the voters went back of their promise to vote for X and 25% of voters went back of their promise to vote for Y and X won by 22 votes. Then, the total number of votes for X are ________.

OpenStudy (anonymous):

3/5th small change

OpenStudy (anonymous):

In these questions the hard part is to construct equations X-Y=22 from the last sentence 3/5 All voted for X but 25% left 25/100=1/4 3/5*1/4=3/20 or 15% of all On the other hand 2/5 voted for Y and 15%left 2/5*3/20=6/100 6 % of all Lets sum up this: 60% voted for X but 15% left 40% voted for Y but 6%left 60-15+6=51% 40-6+15=49% 51% of all votes - 49% of all votes are equal 22 2%of all votes are 22 people All votes given 22*50=1100 51%of 1100=561

OpenStudy (anonymous):

561 is total votes

OpenStudy (anonymous):

Then, the total number of votes for X are ________

OpenStudy (anonymous):

51

OpenStudy (anonymous):

1100 are the total votes

OpenStudy (anonymous):

561 is the answer

OpenStudy (anonymous):

no its wrong

OpenStudy (anonymous):

Lets check 2/5 of 1100= 440 3/5 of 1100= 660 25% of 660= 660/4=165 15% of 440= 66 440+165-66= 539 660+66-165= 561

OpenStudy (anonymous):

And 561-539=22 So I am correct

OpenStudy (anonymous):

something went wrong

OpenStudy (anonymous):

total 100

OpenStudy (anonymous):

You are just throwing numbers now.

OpenStudy (anonymous):

x=60*.85=51 y=40*.75=30

OpenStudy (anonymous):

51+30=81 what happened with the 19 other voters? Poor voters couldnt make it to the booth :(

OpenStudy (anonymous):

x=51+10=61

OpenStudy (anonymous):

y=30+9

OpenStudy (anonymous):

x=61,y=39

OpenStudy (anonymous):

60 voted for X 15 left 40 voted for Y 6 left 60-15+6=51 40-6+15=49 Difference is 2

OpenStudy (anonymous):

Again you do not really do maths

OpenStudy (anonymous):

Just throwing in some random numbers you should leave to politicians.

OpenStudy (anonymous):

then x-y=22,in ur answer it is x-y=2

OpenStudy (anonymous):

In your answer it is 2

OpenStudy (anonymous):

u said difference is 2

OpenStudy (anonymous):

x=61,y=39 that is the number you throw in without any maths. So that means total voters =100 According to the question: 60 voted for X 15 left 40 voted for Y 6 left 60-15+6=51 40-6+15=49 Difference is 2

OpenStudy (anonymous):

2%of all votes are 22 people All votes given 22*50=1100 51%of 1100=561.how 2% of all votes is 22.then how u multiplied 50 with 22 i cant understand

OpenStudy (anonymous):

2%=2/100=1/50 so 100 % is 50*2%